Operations Report – April 5th

Crew 279 Operations Report 05-04-2023

SOL: 3

Name of person filing report: Ioana Dimitrova

Non-nominal systems: robotic observatory and suit number 8

Notes on non-nominal systems:

Robotic observatory: awaiting for a solution from the manufacturer

Suit number 8: Problem is being assessed by crew engineer (details in Summary of suits and radios)


Spirit rover used: no

Hours: 213.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: no

Hours: 119.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: no

Hours: 226.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Perseverance rover used: no

Hours: 260.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

General notes on rovers: none

Summary of Hab operations: Music influence on stress experiment was conducted. Planning of all menus for all Sols to quantify correctly.


Water (static tank): 422,6

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: used from 14:00 to 21:00 for experiments and reports

Summary of suits and radios:

Following yesterday’s problems with helmet fogging up on Suit 8, Engineer tested voltage every 20 min, here are the results:


20 min

40 min

60 min

80 min

100 min

120 min


13,6 V


12,89 V

12,7 V

2,93 V

2,1V (left ventilator turned off)

The EVA’s could last 3h, so it is not normal to have this kind of power loss at one hour and 40 min. I think the battery needs changing. The suit will not be used until a solution is found for changing the battery.

Radios charging nominal. Headset number 5 plastic headband got detached from earphone during EVA, the plastic bond broke. Comms still worked, we were still able to communicate. Engineer reattached the two parts with superglue in the afternoon and the headset is now entirely functional.


Heater: Off

Supplemental light: Off

Harvest: none

Summary of ScienceDome operations : 3D printer reconstruction finished. Bacteria was put in liquid environment culture to be able to be exposed to another environment soon.

Dual split: Heat or AC, Off

Summary of RAM operations: 3D printer tested and print of other test parts will be started tonight. Headset number 5 fixed in the RAM.

Summary of any observatory issues: see above for robotic observatory. Solar observatory not used for cloudy skies. The New Mexico observatory was not open so we got black pictures. See observatory report for more details.

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

Is it normal that when you ignite the stove in the Hab kitchen only on one stovetop, the other stovetops also give out a spark?

Copyright © The Mars Society. All rights reserved. | Main Site